### v2

# Aligning Speech to Languages to Enhance Code-switching Speech Recognition

Abstract
Code-switching (CS) refers to the switching of languages within a speech signal and results in language confusion for automatic speech recognition (ASR). To address language confusion, we introduce a novel language alignment loss into ASR training to align acoustic features to pseudo-language labels learned from the ASR decoder. This approach enables frame-level language identification without the need for frame-level language annotations. To further tackle the complex token alternatives for language modeling in bilingual scenarios, we propose to employ large language models via a generative error correction method. A linguistic hint, which is derived from LAL outputs and decoded hypotheses, is introduced to guide the prompting and enhance the LLM-based generative error correction for CS-ASR. The proposed methods are evaluated on the SEAME dataset and data from the ASRU 2019 Mandarin-English code-switching speech recognition challenge. The incorporation of the proposed language alignment loss improves the CS-ASR performance for both hybrid CTC/attention and Whisper models on both datasets, with only a negligible increase in the number of parameters. This work also highlights the efficacy of language alignment loss in balancing primary-language-dominant bilingual data during training, with an 8.6% relative improvement on the ASRU dataset compared to the baseline model. Performance evaluation using large language models reveals the advantage of the linguistic hint by achieving 14.1% and 5.5% relative improvement on test sets of the ASRU and SEAME datasets, respectively.
Index Terms:
code-switching, speech recognition, alignment, language, large language model
I Introduction
Code-Switch (CS) refers to the switching of languages within a spontaneous multilingual recording. Intra-sentence code-switching occurs when the language changes within a single sentence, while inter-sentence code-switching involves the switching of languages at the sentence boundaries. Unlike monolingual speech, code-switched speech presents a greater challenge for automatic speech recognition (ASR) due to language confusion and the lack of annotated data.
Although a CS-ASR system can function like monolingual ASR by combining language-specific vocabularies [1, 2], recent works address language confusion by incorporating language information. One direct approach is to optimize the ASR and language identification (LID) or diarization (LD) tasks jointly [3, 4, 5, 6, 7]. Here, models learn language information through backpropagation from the LID or LD branch during training [6, 8], while only the ASR output is computed during inference. Beyond joint optimization, approaches based on the bi-encoder and the mixture-of-experts method building upon the Transformer architecture [9, 10, 11, 12], where the models incorporate two encoders pre-trained on monolingual data independently to capture language-specific information. Language-specific modules have also been adopted in other architectures due to their effectiveness in distinguishing languages [13, 14]. In contrast to language-specific encoder modules, the language-aware decoder module has been explored to reduce multilingual contextual information via a language-specific self-attention mechanism within the Transformer decoders [15]. In addition, a conditional factorization method factorizes CS-ASR into two monolingual recognition processes before integrating multiple recognized monolingual segments into a single bilingual sequence [16]. As an extension, a conditionally factorized connectionist temporal classification (CTC) module that allows for a zero-shot setting has been proposed [17, 18]. A factored language model integrating syntactic and semantic features in a code-switched language model has also been explored to enhance CS-ASR [19].
Although existing works achieved reasonable CS-ASR performance, constraints that impede model performance persist. Multilingual ASR systems often benefit from utterance-level one-hot language vectors, as each utterance typically contains only one language [20, 21]. In contrast, code-switching speech signals frequently contain two languages. Therefore, detecting language change points at high granularity (in particular at frame level) is preferable to utterance-level vectors [7]. Despite this need, ground-truth code-switching timestamps are often excluded in code-switching corpora since the human annotation process is resource-intensive and requires expertise in bilingualism. While force alignment offers a potential solution to generate high-granularity timestamps, code-switching speech poses greater challenges for this approach compared to monolingual speech due to increased language confusion [22, 23, 24]. Consequently, language labels generated through force alignment might not be desirable in terms of accuracy and computational cost. These limitations render the application of supervised language identification for joint optimization impractical [5]. Additionally, the accent of the primary language may bias the secondary language in bilingual code-switching speech, making the two languages auditorily similar [24, 25]. Therefore, achieving language identification in accented speech remains challenges [26].
In this paper, we propose to leverage the speech-to-language alignment to improve CS-ASR performance. Central to our approach is introducing a language alignment loss (LAL), which enriches the CS-ASR model with language information without requiring additional annotations such as human labeling or force alignment. The LAL achieves this by utilizing low-granularity yet accurate token-level language information to explicitly guide high-granularity frame-level acoustic features. Unlike auxiliary frame-level LID methods, the proposed approach accounts for the proportion of tokens from each language and their auditory similarity, ensuring a more nuanced alignment between language and acoustic representations. Additionally, as the by-product of LAL, frame-level language predictions can be summarized as an utterance-level hint. This linguistic hint further utilizes the language information derived from speech to facilitate the incorporation of an external large-scale language model (LLM) in ASR via a generative error correction method [27, 28, 29].
The remainder of this paper is organized as follows: Section II introduces the hybrid CTC/attention model, which serves as a baseline model on which we apply the proposed method. The proposed LAL and the incorporation of the code-switching hint in external language modeling are presented in Sections III and III-D, respectively. Information pertaining to the datasets, model configurations, and experiment setup are provided in Section IV. We present results and analysis in Section V before highlighting the advantages and limitations of our proposed methods in Section VI. Finally, we conclude our work in Section VII.
II Preliminary
II-A Conformer-based hybrid CTC/attention ASR model
We employ a Conformer-based hybrid CTC/attention ASR model as the baseline in this paper. The hybrid CTC/attention model comprises an encoder, a decoder, and a CTC module, where the decoder and CTC modules share the encoder outputs [30]. The encoder and decoder modules comprise Conformer encoder and Transformer decoder layers [31, 9], respectively [2].
Given a speech signal, we define its acoustic features as $\mathbf{X}=(\mathbf{x}_{t}\in\mathbb{R}^{F}|t=1,\ldots,T)$ and paired token sequence as $W=(w_{n}\in\mathcal{V}|n=1,\ldots,N)$, where $\mathcal{V}$ is the vocabulary, $T$ and $N$ are the lengths of feature and token sequences, respectively, and $F$ is the dimension of acoustic feature. The encoder generates hidden outputs $\mathbf{H}=(\mathbf{h}_{t}\in\mathbb{R}^{D}|t=1,\ldots,T^{\prime})$ from $\mathbf{X}$, which are subsequently used as inputs into the decoder and CTC modules. Here, $T^{\prime}$ is the length of the hidden output sequence, where $T^{\prime}<T$ due to the subsampling, and $D$ is the dimension of the hidden output. With $\mathbf{H}$, the CTC module computes the token sequences in accordance with the Bayesian decision theory by factorizing $p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)$ as [18]
$p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)$ $=\sum_{Z}p\left(W|Z,\mathbf{X}\right)p\left(Z|\mathbf{X}\right)$ (1)
$\approx\sum_{Z}p\left(W|Z\right)p\left(Z|\mathbf{X}\right),$
where $Z=(z_{t}\in\mathcal{V}\cup\left\{<blank>\right\}|t=1,\ldots,T^{\prime})$ is a framewise token sequence conditioned on $\mathbf{X}$. The variable
$p\left(Z|\mathbf{X}\right)$ $=\prod_{t=1}^{T}p\left(z_{t}|z_{1},\ldots,z_{t-1},\mathbf{X}\right)\approx\prod_{t=1}^{T}p\left(z_{t}|\mathbf{X}\right)$ (2)
is the acoustic model of CTC, where the probabilistic chain rule and the conditional independence assumption have been invoked. Exploiting the Bayes’ rule, the probabilistic chain rule, and the conditional independence assumption, the CTC token model is given by
$p\left(W|Z\right)$ $=\frac{p\left(W\right)p\left(Z|W\right)}{p\left(Z\right)}$ (3)
$=\prod_{t=1}^{T}p\left(z_{t}|z_{1},\ldots,z_{t-1},W\right)\frac{p\left(W\right)}{p\left(Z\right)}$
$\approx\prod_{t=1}^{T}p\left(z_{t}|z_{t-1},W\right)\frac{p\left(W\right)}{p\left(Z\right)}.$
Consequently, the term $p\left(W\right)/p\left(Z\right)$ is often excluded and hence (1) can be rewritten as
$p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)\approx\sum_{Z}\prod_{t=1}^{T}p\left(z_{t}|z_{t-1},W\right)p\left(z_{t}|\mathbf{X}\right).$ (4)
We note that tokens $W$ are embedded into $\mathbf{W}=(\mathbf{w}_{n}\in\mathbb{R}^{D}|n=1,\ldots,N)$ before being fed into the decoder module along with $\mathbf{H}$. The decoder then predicts the next token $w_{n}$ based on historical tokens $w_{1:n-1}$ and $\mathbf{H}$ via
$p\left(w_{n}|w_{1:n-1},\mathbf{X}\right)=\mathrm{Decoder}\left(\mathbf{w}_{1:n-1},\mathbf{H}\right),$ (5)
where $p(w_{n}|w_{1:n-1},\mathbf{X})$ is the posterior of $w_{n}$ given acoustic features and historical tokens, and $\mathrm{Decoder}(\cdot)$ denotes the Transformer decoder. The encoder-decoder module computes the token sequences by factorizing $p_{\mathrm{att}}\left(W|\mathbf{X}\right)$ as
$p_{\mathrm{att}}\left(W|\mathbf{X}\right)\approx\prod_{n=1}^{N}p\left(w_{n}|w_{1:n-1},\mathbf{X}\right).$ (6)
The model is optimized via a multi-task objective function
$\mathcal{L}_{\mathrm{asr}}=\alpha\mathcal{L}_{\mathrm{ctc}}+\left(1-\alpha\right)\mathcal{L}_{\mathrm{att}},$ (7)
where $\mathcal{L}_{\mathrm{ctc}}$ denotes the CTC loss, $\mathcal{L}_{\mathrm{att}}$ denotes the cross-entropy loss with label smoothing for the encoder-decoder branch [32], and $\alpha$ is a parameter associated with multi-task learning. The decoding process aims to maximize the linear combination of the logarithmic CTC and attention objectives such that the decoded token sequence is given by
$\small\widehat{W}=\operatorname*{argmax}_{W}\big{\{}\alpha\mathrm{log}p_{\mathrm{ctc}}\left(W|\mathbf{X}\right)+\left(1-\alpha\right)\mathrm{log}p_{\mathrm{att}}\left(W|\mathbf{X}\right)\big{\}}.$ (8)
Here, $\widehat{W}$ is also referred to as a single hypothesis, and the final hypothesis of the given speech signal is chosen as the one with the highest likelihood among multiple hypotheses generated during the beam search. While the hybrid CTC/attention model has shown to be effective for the CS-ASR task, it performs CS-ASR similarly to monolingual ASR without exploiting any code-switching information, which, as a consequence, limits the CS-ASR performance.
Figure 1: Generative error correction method for improving ASR with an external LLM proposed in [29]. Variable $r$ is the low intrinsic rank, and matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized by a random Gaussian distribution and zero, respectively. The red and black tokens within the N-best hypotheses denote the wrongly and correctly predicted tokens, respectively.
II-B Improving ASR via LLM and efficient fine-tuning
Due to the scarcity of code-switching data, general code-switching language models moderately improve the performance of a CS-ASR system when being incorporated through shallow fusion [4, 7]. Although developing an external language model on monolingual data or synthesized code-switching data has shown to be effective [33, 34, 35], the performance is limited by the domain mismatch. Since large language models have achieved success in natural language processing and have been extended for computer vision and speech signal processing applications [36, 27, 37], we propose to adopt open-source LLMs, which is robust against diverse domains due to the large-scale training data, to improve CS-ASR by addressing the complex token alternatives in bilingual scenarios.
Recent works have also attempted to improve speech recognition through the use of LLMs. A direct method involves prompting an LLM using paired discrete speech and text embeddings [38]. A generative error correction method has also been applied to LLMs [29] as presented in Fig. 1, where the final prediction is generated by summarizing and correcting the N-best ASR hypotheses [39]. This approach has shown to be effective in monolingual ASR. However, code-switching leads to more token alternatives that are of similar auditory or syntactic characteristics compared to a monolingual application—direct transference of the generative error correction to CS-ASR may not be desirable.
In addition, to fine-tune an LLM efficiently, low-rank adaptation (LoRA) [40] has been proposed. As shown in Fig. 1, computational complexity is reduced by freezing the pre-trained LLM and injecting trainable rank decomposition matrices $\mathbf{A}$ and $\mathbf{B}$ into its Transformer-based layers. The forward pass is then defined as the linear combination of the pre-trained model $\mathbf{M}_{0}$ and the trained decomposed matrices $\mathbf{A}$ and $\mathbf{B}$ such that
$\left(\mathbf{M}_{0}+\mathbf{\Delta M}\right)\mathbf{X}=\left(\mathbf{M}_{0}+\mathbf{AB}\right)\mathbf{X},$ (9)
where $\mathbf{\Delta M}$ is the model parameters of the model update. Matrices $\mathbf{A}$ and $\mathbf{B}$ are initialized as by random Gaussian and zero values, respectively, so that $\mathbf{\Delta M}=\mathbf{AB}=0$ at the beginning of training.
III Language Alignment Loss
Since multilingual ASR benefits from the supplementary language information offered by utterance-level one-hot language vectors, we propose to enhance CS-ASR performance by incorporating frame-level language information, enabling the detection of code-switching at high granularity. In this work, we introduce the LAL that is incorporated to the encoder-decoder framework to capture language information as shown in Fig. 2.
III-A Frame-level language identification
To capture frame-level language information, we employ a linear layer as a built-in language classifier. This layer takes the hidden output unit of the encoder module as its input and aims to generate a language decision for each hidden output unit.
Due to the lack of frame-level ground-truth or gold-standard language labels, existing works usually perform language identification in an unsupervised manner [11, 7]. Nevertheless, this unsupervised frame-level classification extends beyond language identities and includes elements such as phonemic or domain information. Therefore, incorporating language information becomes important in guiding the unsupervised language identification process.
Although frame-level ground-truth language timestamps are unavailable, token-level language information can readily be inferred from text, particularly in cases where there is a notable contrast in character structure or morphology between the two languages. Past research has investigated the conversion of byte-pair encoding (BPE) tokens into their respective language labels to facilitate language identification or diarization [41, 4, 42]. Here, we employ a similar conversion strategy, where sub-tokens like BPE tokens are first transformed into token-level labels corresponding to their respective languages. The pseudo-frame-level language annotations are subsequently extracted by aligning frames to these token-level language labels.
Figure 2: The hybrid CTC/attention model (in blue) with language alignment process.
III-B Aligning frames with token-level language labels
The encoder-decoder model achieves ASR by mapping speech features to tokens. The alignment between speech frames (i.e., encoder hidden outputs) and text (i.e., tokens) is inherently learned through the cross-attention process illustrated on the left of Fig. 3. Given multiple attention heads within the last Transformer decoder layers, their cross-attention matrices are averaged to generate the weight matrix $\mathbf{Atten}\in\mathbb{R}^{T^{\prime}\times N}$ of the speech-to-text alignment shown in the top matrix of Fig. 3. The alignment between hidden outputs and languages can then be derived from the frame-to-token weight matrix illustrated by the transition from the top to bottom matrices in Fig 3. Specifically, the averaged cross-attention weight matrix can be decomposed into vectors along $T^{\prime}$, being $\mathbf{Atten}=(\mathbf{atten}_{t}\in\mathbb{R}^{N}|t=1,\ldots,T^{\prime})$. Each element in $\mathbf{atten}_{t}$ denotes the attention weight corresponding to a BPE token within the input token sequence. Each frame can then be assigned a pseudo language label which is the language of the BPE token corresponding to the highest weight. This is achieved via
$\mathbf{y}_{t}=\mathrm{T2L}\big{(}\operatorname*{argmax}_{n}\left(\mathbf{atten}_{t}\right)\big{)},$ (10)
[…]
The pseudo language labels are concatenated to form a sequence $\mathbf{Y}=(\mathbf{y}_{t}\in\mathbb{R}^{C}|t=1,\ldots,T^{\prime})$. Here, $T^{\prime}$ is the number of hidden output units and $C$ is the number of languages, where all special tokens are treated as a third language in addition to the two target languages. Each hidden output unit is subsequently projected to a language decision $\mathbf{y}_{t}$ by a language classifier comprising one linear layer
$\mathbf{\widehat{y}}_{t}=\mathrm{Linear}\left(\mathbf{h}_{t}\right),$ (11)
where $\mathrm{Linear}\left(\cdot\right)$ denotes computations within a linear layer. Defining $\mathrm{exp}(\cdot)$ as the exponential operation, the language alignment loss for each speech sample is computed via a cross-entropy function
$\mathcal{L}_{\mathrm{lid}}=\frac{-1}{T^{\prime}}\sum_{t=1}^{T^{\prime}}\sum_{c=1}^{C}\mathrm{log}\frac{\mathrm{exp}\left(\mathbf{\widehat{y}}_{t,c}\right)}{\sum_{i=1}^{C}\mathrm{exp}\left(\mathbf{\widehat{y}}_{t,i}\right)}\mathbf{y}_{t,c},$ (12)
where $\mathbf{\widehat{y}}_{t,i}$ denotes the value at the $i$th dimension of the language prediction vector for the $t$th hidden output vector. Similarly, $\mathbf{y}_{t,c}$ denotes the value at the $c$th dimension of the one-hot pseudo language label vector for the $t$th hidden output vector such that $\mathbf{y}_{t,c}=1$ for the target language with the remaining elements being zero.
The key distinction between the proposed LAL and general LID lies in the ability of the proposed LAL to iteratively refine pseudo-labels during training, enabling a mutual optimization process between ASR and LID tasks. The performance of the CS-ASR model improves during model training as the language information is incorporated, reaching its peak upon being optimized completely. The heightened ASR performance contributes to increasingly accurate language labels. This improvement, in turn, leads to higher language identification performance. Therefore, extracting pseudo language labels and language identification has become integral to the iterative optimization process during training.
Figure 3: The pseudo frame-to-language label conversion process. On the right side, the upper matrix represents the cross-attention weight matrix averaged over all attention heads within the last decoder layer. The middle matrix represents the converted frame-to-language weight matrix and the pseudo language labels are illustrated at the bottom. In this illustrative example, red and green cells denote the English and Mandarin tokens, respectively.
III-C Balancing training via toke-level language weights
The grammatical structure of code-switching data defines the matrix language as the main language and the embedded language as the secondary language [43]. While code-switching corpora such as SEAME do not exhibit a dominant language [24, 19], the matrix language may prevail in some code-switching corpora, resulting in an imbalanced distribution of tokens and speech frames between the two languages [25]. Therefore, a CS-ASR model trained on such imbalanced data may overfit the matrix language and underfit the embedded language.
While computing the number of speech frames for each language is challenging, the token distribution within a code-switching corpus can be assessed before training. Therefore, we propose to address the imbalance issue by incorporating language weights in (12), being a weighted cross-entropy function as
$\mathcal{L}_{\mathrm{lal}}=\frac{-1}{T^{\prime}}\sum_{t=1}^{T^{\prime}}\sum_{c=1}^{C}w^{\mathrm{lang}}_{c}\mathrm{log}\frac{\mathrm{exp}\left(\mathbf{\widehat{y}}_{t,c}\right)}{\sum_{i=1}^{C}\mathrm{exp}\left(\mathbf{\widehat{y}}_{t,i}\right)}\mathbf{y}_{t,c},$ (13)
where $w^{\mathrm{lang}}_{c}$ denotes the normalized language weight of language $c$ and can be made inversely proportional to its token count in the training data. However, the token ratio may not align precisely with speech frames due to variations in speech rates across languages. We therefore propose to tune them initially based on the token ratio and accounting for speech rate variations when applying language weights to the LAL. The relationship between the initial language weights can be illustrated by
$\frac{w^{\mathrm{lang}}_{\mathrm{Eng}}}{w^{\mathrm{lang}}_{\mathrm{Man}}}\propto\frac{\mathrm{Count}(\mathrm{tokens}_{\mathrm{Man}})}{\mathrm{Count}(\mathrm{tokens}_{\mathrm{Eng}})},$ (14)
where $\mathrm{Count}(\cdot)$ denotes the number of BPE tokens belonging to the language. In particular, fewer tokens in the training data and a higher speech rate result in a higher language weight.
With the above, the CS-ASR model is optimized via an objective function being the sum of CTC loss, encode-decoder loss, and the proposed LAL such that
$\mathcal{L}_{\mathrm{asr}}=\alpha\mathcal{L}_{\mathrm{ctc}}+\left(1-\alpha\right)\mathcal{L}_{\mathrm{att}}+\beta\mathcal{L}_{\mathrm{lal}},$ (15)
where $\beta$ denotes the weight of the language alignment loss during training. The decoding process is similar to (8).
III-D Linguistic hint for prompting LLM
With reference to Fig. 1, we propose to adopt the LLM-based generative error correction method to improve CS-ASR [29]. To fine-tune the LLM efficiently, LoRA is employed while the LLM is kept frozen during training as illustrated in Section II-B. The prompt originally designed in [29] is used as shown in Fig. 4(a), where the N-best hypotheses are extracted from the ASR output before being inserted in the prompt. The LLM is then optimized to deduce the correct transcript, which is the ground-truth transcription shown in Fig. 4 during training, by leveraging the information provided in these hypotheses. However, code-switching gives rise to more intricate token alternatives with similar auditory or syntactic characteristics compared to a monolingual application. This complexity persists when performing generative error correction for ASR hypotheses. Inspired by the use of chain-of-thought as additional supervision when fine-tuning LLM for downstream tasks [44, 28], we propose to employ an additional linguistic hint during prompting to address the aforementioned challenge in CS-ASR.
To this end, various methods for linguistic hint extraction can be used. An acoustic-biased linguistic hint can be derived from the by-product of the proposed LAL. As shown in Fig. 2, frame-level language predictions are first normalized using the softmax function before generating an utterance-level language decision. In the context of this work, this decision can either be monolingual or multilingual, with the former providing a single language code. In addition, the linguistic hint can be obtained from the decoded hypotheses (i.e., text information), allowing it to be employed in conjunction with the acoustic-biased hint through the weighted voting mechanism.
The linguistic hint is then inserted in the used prompt during fine-tuning as shown in Fig. 4(b). We propose two types of hints—the monolingual hint, where only $<$language id$>$ words are included in the transcription, and the code-switching hint, where speech is multilingual and that $<$both language ids$>$ words are included in the transcription.
Figure 4: The original (a) and proposed (b) prompts during fine-tuning LLM and performing generative error correction, language id is selected from two languages (“English” and “Mandarin” in this work) and “multilingual”.
IV Dataset, experiment, and model configuration
IV-A Datasets
We conducted the experiments on data extracted from the ASRU 2019 Mandarin-English code-switching speech recognition challenge and the SEAME dataset [24, 25]. The ASRU 2019 Mandarin-English code-switching speech recognition challenge consists of four datasets, including a 500-hour Mandarin-only training set, a 200-hour intra-sentence English-Mandarin code-switching training set, a 40-hour intra-sentence English-Mandarin code-switching development set, and a 20-hour intra-sentence English-Mandarin code-switching test set. In the experiments, the models were trained on the 200-hour CS training set, validated on the development set, and evaluated on the test set. The SEAME dataset, on the other hand, is a Mandarin-English code-switching corpus containing spontaneous conversational speech [24]. This dataset encompasses both intra- and inter-sentence code-switching speech. We divided the SEAME dataset into a 96.6-hour training set, a 4.9-hour validation set, and two test sets being denoted as $\text{test}_{\texttt{man}}$ and $\text{test}_{\texttt{sge}}$ following the same partitioning method described in [3]. Detailed information about the test sets is provided in Table I, while Table II highlights the duration ratio of each language, where the language labels are annotated at the utterance level.
While both datasets involve English-Mandarin code-switching, the primary distinction between them lies in the accent. The ASRU dataset, recorded in mainland China, is characterized by a dominant Chinese accent and text. In the training and development sets, each sentence, on average, consists of 8.6 Chinese characters and 1.6 English words. In contrast, the SEAME dataset comprises audio recordings from Singapore and Malaysia, featuring South-East Asian accents. Additionally, code-switching occurs more frequently within the SEAME dataset compared to the ASRU data due to the bilingual education and language policies in Singapore and Malaysia [45]. This suggests that the SEAME data might pose greater challenges for CS-ASR compared to the ASRU data.
TABLE I: Details of two datasets in terms of division and durations
      Corpus       Subset       Duration (hours)
      ASRU       train       193.0
      dev       21.3
      test       20.4
      SEAME       train       96.6
      dev       4.9
      $\text{test}_{\texttt{man}}$       7.5
      $\text{test}_{\texttt{sge}}$       3.9
TABLE II: Duration ratio for Mandarin, English, and code-switching speech of test sets in ASRU and SEAME data. The language annotations are at utterance level
Subset Duration ratio (%)
Mandarin English Code-Switching
ASRU test 0 0 100
$\text{SEAME test}_{\texttt{man}}$ 14 7 79
$\text{SEAME test}_{\texttt{sge}}$ 6 41 53
IV-B Data preprocessing
Since the SEAME dataset contains a small training set of about 98 hrs, we augmented the training data using speed perturbation and SpecAugment [46, 47] for training models from scratch. Two training strategies were adopted, where one develops the model on data without speed perturbation while the other trains the model on the entire augmented data. The speech perturbation was applied with factors 0.9, 1.0, and 1.1. With respect to the ASRU dataset, only SpecAugment was applied for data augmentation. SpecAugment adopted the default setup in ESPnet for two datasets [2]. The time-warp mask size was set to five, and two time and frequency masks were applied, with their lengths uniformly selected from the range of 0 to 40 for time masks and 0 to 30 for frequency masks. Speech samples within both corpora were segmented into durations ranging from 0.1 to 20 s. We extracted $F=80$ dimensional log-Mel-Fbank features for each speech segment before applying the cepstral mean and variance normalization.
For the train-from-scratch models, we employed BPE to tokenize the English words in the two English-Mandarin code-switching corpora and split all Mandarin words into individual characters. For the SEAME dataset, this resulted in a total of $V=5,628$ tokens, including 3,000 English BPE tokens, 2,624 Mandarin characters, and four special tokens for $<$unk$>$, $<$noise$>$, $<$blank$>$, and $<$sos/eos$>$. For the ASRU data, the same tokenization process yielded a total of $V=6,923$ tokens, including 3,000 English BPE tokens, 3,920 Mandarin characters, and three special tokens for $<$unk$>$, $<$blank$>$, and $<$sos/eos$>$. The variable $C=3$ denotes three language classes converted from special, English, and Mandarin tokens.
We fine-tuned the Chinese LLaMA-2 on a subset of the SEAME training set comprising approximately 60,000 speech segments and the development set of the ASRU data comprising approximately 20,000 speech segments, respectively. N-best lists of the above speech segments were subsequently extracted and incorporated into the prompts. In addition, we removed the $<$noise$>$ and $<$unk$>$ labels from the N-best list since the LLM can hardly address these special tokens.
For experiments involving Whisper models, we adopted the data preprocessing and tokenization methods outlined in [37]. Force alignment is utilized to generate frame-level language annotations and is performed using the fine-tuned Whisper-small (FA-LB) and Whisper-large (FA-UB) models.
Figure 5: The convergence of the proposed language alignment loss (y-axis for the left figures) during training and validation on the (a) ASRU data and (c) SEAME data, and the corresponding MER (y-axis for the right figures) on the validation set of (b) the ASRU data and (d) SEAME data, against training epochs (x-axis).
IV-C Model configuration and experiment setup
The baseline model is a Conformer-based hybrid CTC/Attention ASR model comprising twelve Conformer encoder layers, six Transformer decoder layers, and a CTC module. The Conformer employs the macaron structure [31], where its convolutional neural network (CNN) module has a kernel size of 15. The swish activation function is applied [48]. A CNN layer first subsamples the input features and projects them into $D=256$ dimensions before feeding into the macaron modules. All attention layers within the encoder and decoder modules have four attention heads with input and output dimensions being $D=256$, and the inner layer of the position-wise feed-forward network is of dimensionality 2048.
The bi-encoder and language posterior bias methods have also been implemented as baselines [11, 4]. In the bi-encoder CS-ASR system, we replicated the encoder module of the baseline model described in the previous paragraph to form the bi-encoder module. The outputs of these two encoders are fed into a linear layer, which serves as a built-in LID module with an output dimensionality of two (corresponding to two languages), before calculating the weighted encoder outputs. It is useful to note that the encoder modules of the bi-encoder method are pre-trained on English and Mandarin corpora separately. Nevertheless, this pre-training was deliberately omitted to ensure a fair comparison. For the language posterior bias-based (LPB) CS-ASR model, the model configuration recommended in [4] was used. The LPB system closely resembles the baseline hybrid CTC/attention model, with an additional language diarization decoder. This language diarization decoder mirrors the structure of the ASR decoder.
We trained the baseline model on the ASRU and SEAME data for seventy and fifty epochs, respectively. The proposed method and reproduced works were trained for an additional ten epochs due to their more complex objective functions or parameters. All models were optimized by an Adam optimizer on two RTX 3090 GPUs. The learning rate increased from 0 to 0.001 over 25,000 update steps, followed by a cosine annealing decay. Parameter $\alpha=0.3$ and a label smoothing factor of 0.1 were used in (7) and (15). The ten best models during validation were averaged for inference. We adopted the ten-best beam search method with $\alpha=0.4$ in (8).
Experiments associated with the proposed linguistic hint utilized a Chinese LLaMA2-7B model  [^1] https://huggingface.co/hfl/chinese-llama-2-7b. The 5-best hypotheses were first extracted from the ASR output before being used in the prompt to fine-tune the LLM. Fine-tuning was performed via LoRA with rank $r=4$ and the AdamW optimizer for ten epochs with a batch size of 128 on the SEAME data, where LoRA is applied to the query and value modules within the self-attention modules. The learning rate was increased from 0 to 0.0002 over 100 update steps, followed by a linear decay. During inference, a temperature of 0.7 was applied to allow for the creativity of the LLM, while other hyper-parameters were fixed at their default settings. We evaluated all methods by employing the MER comprising WER for English and CER for Mandarin.
Experiments relevant to Whisper models utilized the Whisper-small model [37]. Fine-tuning was conducted using an AdamW optimizer over five epochs, with a learning rate increased from 0 to $1\times 10^{-5}$ over 10,000 update steps, followed by a linear decay. The training employed a batch size of eight samples, with gradients accumulated every two updates. Inference was carried out using the best-performing model on the development sets, with a beam size of one. Language prompts $<$zh$>$ and $<$en$>$ are used for the ASRU and SEAME data, respectively. Since Whisper models were pre-trained on large-scale data, $\beta$ is set to 0.05 and 0.01 for ASRU and SEAME data, respectively, to prevent the LAL or frame-level LID from exerting excessive influence during fine-tuning. The frame-to-token alignment for LAL is computed via the cross-attention matrices within the last decoder layer. LoRA was applied to the query, key, value, and linear modules with rank $r=8$ during parameter-efficient fine-tuning.
Evaluation of CS-ASR systems was quantified via the mix error rate (MER) comprising word error rate (WER) for English and character error rate (CER) for Mandarin. LID accuracy was calculated at the utterance level.
V Results and analysis
V-A Impact of $\beta$ values
To assess the impact of $\beta$ on the CS-ASR performance for the proposed LAL, we first compare the performance of the hybrid CTC/attention model optimized with LAL via (15) for various $\beta$ values on the ASRU data. Results summarized in Table III highlight that with the incorporation of LAL during training, the hybrid CTC/attention model consistently outperforms the vanilla model. Notably, the best MER of 11.9% is observed when $\beta=1.5$, achieving a relative improvement of 7.03% compared to the vanilla model. In addition, this performance improvement remains consistent across various LAL parameters, particularly within the range of $1.0\leq\beta\leq 3.0$.
TABLE III: Performance evaluation of the proposed method on the ASRU data with different $\beta$ values in terms of substitutions, deletions, insertions, and the total MER (%)
Method $\beta$ for LAL Sub $\downarrow$ Del $\downarrow$ Ins $\downarrow$ MER $\downarrow$
Hybrid CTC/atten 0 11.5 0.6 0.7 12.8
0.5 11.2 0.6 0.7 12.4
1.0 10.9 0.6 0.7 12.1
1.5 10.7 0.5 0.6 11.9
+ LAL 2.0 10.8 0.5 0.6 12.0
2.5 10.9 0.5 0.6 12.0
3.0 10.9 0.5 0.6 12.0
4.0 11.4 0.6 0.6 12.6
5.0 11.0 0.6 0.6 12.2
We next conduct experiments on the SEAME dataset without using speed perturbation during training and results are presented in Table IV. Similar to the ASRU dataset, the hybrid CTC/attention model that incorporates the proposed LAL outperforms the vanilla configuration. In addition, the hybrid CTC/attention model with LAL exhibits consistently high performance for $1.0\leq\beta\leq 3.0$. This robustness on various datasets and $\beta$ values indicates the effectiveness of the proposed LAL. As opposed to results achieved on the ASRU dataset, the highest overall performance on the SEAME data is achieved when $\beta=0.1,0.5,$ and 3.0. Due to the different duration ratios of languages of the two test sets, the above implies that a lower $\beta$ value leads to a higher performance on the dataset containing predominantly monolingual data.
The impact of $\beta$ values can also be discussed from the perspective of accent. Compared to the ASRU data, which is characterized by a dominant Chinese accent, the SEAME data is characterized by Southeast Asian accents. Due to the impact of the Southeast Asian accents on the two languages, the English and Mandarin speech in the SEAME dataset can exhibit auditory similarity, making LID for the SEAME data more challenging than for the ASRU data. Therefore, a larger $\beta$ value is required to achieve high performance when training the CS-ASR model on the SEAME data.
TABLE IV: Performance evaluation of the proposed method trained without speed perturbation on the SEAME dataset with different $\beta$ values in terms of substitutions, deletions, insertions, and the total MER (%)
Method $\beta$ for LAL Subset Sub $\downarrow$ Del $\downarrow$ Ins $\downarrow$ MER $\downarrow$
Hybrid CTC/atten 0 $\text{test}_{\texttt{man}}$ 12.0 3.0 2.2 17.2
$\text{test}_{\texttt{sge}}$ 17.3 4.1 3.1 24.5
+ LAL 0.1 $\text{test}_{\texttt{man}}$ 11.8 2.9 2.2 16.8
$\text{test}_{\texttt{sge}}$ 17.0 4.0 3.0 23.9
0.5 $\text{test}_{\texttt{man}}$ 11.8 2.9 2.2 16.8
$\text{test}_{\texttt{sge}}$ 17.0 4.0 3.0 23.9
1.0 $\text{test}_{\texttt{man}}$ 11.7 3.0 2.1 16.8
$\text{test}_{\texttt{sge}}$ 17.1 4.0 3.1 24.1
2.0 $\text{test}_{\texttt{man}}$ 11.9 3.0 2.0 16.9
$\text{test}_{\texttt{sge}}$ 17.0 4.0 3.1 24.1
3.0 $\text{test}_{\texttt{man}}$ 11.7 3.0 2.0 16.7
$\text{test}_{\texttt{sge}}$ 17.0 4.0 3.0 24.0
4.0 $\text{test}_{\texttt{man}}$ 11.9 3.0 2.0 16.9
$\text{test}_{\texttt{sge}}$ 17.1 4.2 3.0 24.2
5.0 $\text{test}_{\texttt{man}}$ 11.8 3.0 2.2 17.0
$\text{test}_{\texttt{sge}}$ 17.4 4.1 3.1 24.5
TABLE V: Performance of the proposed method with $\beta=1.5$ and different language weights to balance the Mandarin-dominant ASRU data during training in terms of English WER, Mandarin CER, and MER (%). “$ft$” denotes fine-tuning
Language weights ($\beta=1.5$) ASRU test
Other Eng Man Eng $\downarrow$ Man $\downarrow$ MER $\downarrow$
vanilla (all 1) 35.4 9.3 11.9
1 10 1 35.2 9.3 11.8
1 50 1 35.3 9.3 11.8
1 100 1 35.1 9.2 11.7
1 100 100 35.8 9.5 12.1
1 1000 1 35.4 9.3 11.8
V-B Balancing the ASRU dataset during training
The effect of language weights employed in (13) for balancing the Mandarin-dominant ASRU data is shown in Table V. It is worth noting that, similar to the number of English frames, the number of frames for the class “other” is also significantly lower than that for the Mandarin frames. Since the class “other” does not contribute to the language identities in the CS-ASR task, this class is not balanced during training and the weight is always set to $w_{o}ther^{\mathrm{lang}}=1$ except for the learnable language weight. In this work, $w_{c}^{\mathrm{lang}}=1$ for all classes is the default setup.
Results presented show that the CS-ASR model gains moderate performance improvement on both English and Mandarin data from high English weights, where the weights of “other” and “Mandarin” were set to 1. This is consistent with our assumption that a high English weight can achieve balance for the secondary language during training, which consequently improves the model performance. Moreover, the highest performance is achieved when the weight of English is set to 100. It is also worth noting that the ratio between English to Mandarin weights is higher than the token ratio due to the difference between their speech rates.
TABLE VI: Results of the proposed method on SEAME dataset after employing speed perturbation with factor 0.9, 1.0, and 1.1 during training in terms of substitutions, deletions, insertions, and the total MER (%)
Method $\beta$ for LAL Subset Sub $\downarrow$ Del $\downarrow$ Ins $\downarrow$ MER $\downarrow$
Hybrid CTC/atten 0 $\text{test}_{\texttt{man}}$ 11.5 3.0 2.0 16.6
$\text{test}_{\texttt{sge}}$ 16.4 3.9 3.0 23.3
+ LAL 0.5 $\text{test}_{\texttt{man}}$ 11.6 3.0 2.1 16.7
$\text{test}_{\texttt{sge}}$ 16.4 4.1 3.0 23.5
1.0 $\text{test}_{\texttt{man}}$ 11.6 2.9 2.2 16.7
$\text{test}_{\texttt{sge}}$ 16.4 4.0 3.1 23.6
2.0 $\text{test}_{\texttt{man}}$ 11.5 3.0 2.0 16.5
$\text{test}_{\texttt{sge}}$ 16.5 4.1 3.0 23.5
3.0 $\text{test}_{\texttt{man}}$ 11.3 3.0 2.1 16.4
$\text{test}_{\texttt{sge}}$ 16.2 4.0 2.9 23.3
4.0 $\text{test}_{\texttt{man}}$ 11.7 3.1 2.1 16.8
$\text{test}_{\texttt{sge}}$ 16.6 3.9 3.0 23.6
5.0 $\text{test}_{\texttt{man}}$ 11.6 3.0 2.1 16.7
$\text{test}_{\texttt{sge}}$ 16.6 4.0 3.1 23.6
V-C Impact of speed perturbation
As discussed in Section V-B, the frame-level language identification can be affected by speech rate. We, therefore, investigate the impact of the speed perturbation on the CS-ASR performance when LAL is incorporated. The experiments were conducted on the SEAME dataset and in contrast to the experiments shown in Table IV, we utilize speed perturbation to augment the training data with results being summarized in Table VI.
Results presented show that the hybrid CTC/attention model achieves higher performance after employing speed perturbation as data augmentation. However, due to speech rate variations, one token can be aligned with a different number of acoustic frames when speed perturbation is applied. Therefore, further incorporation of the proposed LAL may lead to lower performance. In addition, the hybrid CTC/attention model with LAL being incorporated achieves the highest performance with $\beta=3$. This is consistent with the performance of its counterpart on the SEAME dataset without speed perturbation. It is useful to note that the proposed method achieves the highest performance on the ASRU dataset when $\beta=1.5$. Since a higher $\beta$ value is required for the SEAME data to achieve the highest performance, this also underpins that the SEAME dataset presents greater challenges in language discrimination than the ASRU dataset.
In addition, when speech perturbation is not used, the results presented in Tables III and IV show that the proposed LAL can consistently improve the CS-ASR performance compared to the baseline system. Similar to other hyper-parameters such as learning rate and $\alpha$ in (7) and (15) which require fine-tuning to achieve optimal performance, a fine-tuned $\beta$ results in significant improvements over the baseline model.
TABLE VII: Performance evaluation of the proposed method and state-of-the-art approaches on test data from the ASRU 2019 challenge and SEAME dataset in terms of English WER, Mandarin CER, and MER (%). “PT” denotes the model is pre-trained, “$ft$” denotes fine-tuning, frame-level labels for “LID” are generated via force alignment, “FA-LB” and “FA-UB” denote the force alignment computed via fine-tuned Whisper-small (lower bound) and Whisper-large (upper bound) models, respectively
Method #Train. PT ASRU test $\text{SEMAE test}_{\texttt{man}}$ $\text{SEAME test}_{\texttt{sge}}$
[…]
Hybrid CTC/atten 48.27 M 37.1 10.2 12.8 29.2 15.0 16.6 28.2 22.0 23.3
+ LPB [4] 79.90 M 35.3 9.22 11.8 - - 16.3 - - 22.9
+ LAL (ours) 48.27 M 35.1 9.18 11.7 29.1 14.8 16.4 28.3 21.7 23.3
Bi-encoder [11] 59.98 M 36.0 9.81 12.4 29.2 15.0 16.5 28.2 21.7 23.2
Whisper-small [37] - - - 24.9 - - 90.8 - - 69.7
+ LoRA $ft$ 3.24 M - - 10.9 - - 16.0 - - 21.2
+ $ft$ 244.59 M - - 10.1 - - 14.3 - - 20.0
+ $ft$ w/ prompt [49] 244.59 M - - 10.1 - - 15.1 - - 20.9
+ $ft$ w/ LID FA-LB 244.59 M - - 9.91 - - 14.7 - - 19.9
+ $ft$ w/ LID FA-UB 244.59 M - - 9.61 - - 14.3 - - 19.3
+ $ft$ w/ LAL (ours) 244.59 M - - 9.78 - - 14.5 - - 19.5
V-D Comparison with state-of-the-art methods
We compared the performance of the proposed method against state-of-the-art approaches, including train-from-scratch and pre-trained models. These methods are evaluated on both model parameters and the CS-ASR performance is outlined in Tables VII. For the ASRU dataset, the hybrid CTC/attention model incorporating LAL with language weights achieves the best MER of 11.7% among the train-from-scratch approaches under consideration. While the bi-encoder and LPB methods exhibit higher performance compared to the baseline model, these approaches require an additional Conformer encoder and a Transformer decoder, respectively, resulting in a notable increase in model parameters.
On the SEAME dataset, the hybrid CTC/attention model with the proposed LAL demonstrates a modest performance improvement compared to the vanilla model. Although the LPB method outperforms the proposed method, it requires a higher number of model parameters. In light of the ASRU dataset results, the above findings suggest that integrating the proposed LAL is a lightweight and efficient approach to attain high performance compared to other approaches.
The proposed method has also been validated through fine-tuning the Whisper model. The Whisper model initially demonstrates low zero-shot performance on both datasets, with particularly high error rates of 90.8% and 69.7% observed on the SEAME test sets. However, models fine-tuned on the respective training sets demonstrate improved performance over the pre-trained Whisper, outperforming models trained from scratch. Fine-tuning with the proposed LAL leads to a higher performance compared to pure fine-tuning. This is consistent with the results observed in the hybrid CTC/attention model, highlighting the effectiveness of the proposed method.
Fine-tuning Whisper with either LID or LAL consistently results in improved performance compared to standard fine-tuning, indicating that incorporating auxiliary frame-level LID enhances CS-ASR performance. Furthermore, the Whisper model fine-tuned with the proposed LAL outperforms models that rely on auxiliary LID using language labels derived from FA-LB. This observation supports our statement in Section III-B that the proposed LAL integrates the CS-ASR and LID into a mutual optimization process during training, leading to better performance than conventional LID approaches. However, the proposed LAL shows lower performance than jointly optimizing ASR and LID with FA-UB. This implies that the proposed LAL generates better pseudo labels compared to force alignment computed by models with comparable performance levels, though these labels may still fall short of the accuracy provided by highly precise language labels, such as force alignment computed by fine-tuned Whisper-large model and gold-standard annotations.
TABLE VIII: Performance comparison of different types of linguistic hint used for prompting LLM via generative error correction by employing MER (%) and utterance-level language identification accuracy (Acc %) for English, Mandarin, and Code-switching. The N-best list is decoded from the hybrid CTC/attention model with LAL
Ling. hint type $\text{test}_{\texttt{man}}$ $\text{test}_{\texttt{sge}}$ ASRU test
MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$
no hint 16.4 - 23.3 - 11.0 -
LID output (LAL) 17.0 80.3 24.4 84.3 11.0 95.2
1st hypo. 16.5 93.3 23.1 92.5 11.0 99.7
hypos. vote 16.6 93.2 23.0 93.0 11.0 99.8
LAL & hypos. vote 16.6 92.9 23.1 92.5 11.0 99.9
groundtruth 15.7 - 22.0 - 11.0 -
TABLE IX: Performance evaluation of prompting LLM with generative error correction after incorporating the linguistic hint by employing MER (%). “$ft$” denotes fine-tuning, “$gt$” denotes the linguistic hint using ground-truth utterance-level language label, and “$pred$” denotes the predicted linguistic hint
Method SEAME ASRU
$\text{test}_{\texttt{man}}$ $\text{test}_{\texttt{sge}}$ test
Hybrid CTC/atten 16.6 23.3 12.8
+ LM $latefusion$ 16.4 23.0 12.6
Hybrid CTC/atten w/ LAL 16.4 23.3 11.7
+ LM $latefusion$ 16.4 23.1 11.9
+ LLM LoRA $ft$ 16.4 23.3 11.0
+ LLM LoRA $ft$ + ling. hint $(pred)$ 16.5 23.1 11.0
+ LLM LoRA $ft$ + ling. hint $(gt)$ 15.7 22.0 11.0
V-E Incorporating LLM with linguistic hint
To further enhance the CS-ASR performance, we integrate the LLM through prompting to perform generative error correction on the decoded N-best list. We compare the performance of the prompt in terms of MER and utterance-level LID accuracy before and after incorporating our proposed linguistic hints. The results on SEAME $\text{test}_{\texttt{man}}$, $\text{test}_{\texttt{man}}$, and the ASRU test set are presented in Table VIII. Here, the LID accuracy for the hypothesis within the N-best list is determined by summarizing the languages of tokens and that “LID output” denotes the by-product of the proposed LAL.
Similar to existing works [50, 8], it is not surprising that the LID accuracy achieved with the incorporation of textual information is higher than the by-product of the proposed LAL. This can be attributed to the fact that an ASR system models both acoustic and language characteristics, whereas an LID system generally focuses solely on acoustic information. Furthermore, a decrease in LID accuracy often results in a corresponding degradation in CS-ASR performance of the generated linguistic hints. The prompt incorporating the linguistic hint, which serves as the ground-truth language label, therefore exhibits significantly higher performance on the SEAME and ASRU datasets compared to other prompts.
Notwithstanding the above, prompts with other linguistic hints achieve comparable performance to the prompt without the hint, albeit with a moderately lower overall performance in terms of MER. This observation suggests that linguistic hints may introduce a substantial bias into the generative error correction process. In particular, the utilization of a correct linguistic hint in a prompt enhances CS-ASR performance. Conversely, the misclassification of a linguistic hint results in an increased error rate. This implies that the proposed linguistic hint can potentially improve CS-ASR performance, particularly when an accurate language label is available.
The performance of CS-ASR systems with external language modeling are summarized in Table IX. We observe that incorporating an LM via shallow fusion improves the CS-ASR performance. The use of LLM does not lead to performance improvement on the SEAME dataset unless accompanied by the ground-truth linguistic hint. In contrast, the LLM can improve the performance on the ASRU dataset, achieving an MER of 11.0% for all types of prompts. We analyzed the performance by considering the data distribution as shown in Table II. Since all utterances in the ASRU test set are code-switched, these predicted linguistic hints all exhibit high LID accuracy, and thus show comparable CS-ASR performance.
Since the LLM-based generative error correction method improves the CS-ASR performance over the baseline system on the ASRU data, results on the SEAME data imply that this performance mismatch may be due to differences between the two datasets. As illustrated in Section IV-A, the SEAME dataset comprises spontaneous speech with frequent code-switching and approximately balanced English and Mandarin while the ASRU dataset consists of mostly read speech and Mandarin-dominant text. Since we employ a Chinese LLaMA model to perform general error correction, text data within the SEAME dataset can be significantly mismatched with the training data of Chinese LLaMA. This mismatch leads to a low performance of the LLM-based generative error correction method on the SEAME data. Moreover, incorporating the ground-truth linguist hint in the prompt results in significantly higher performance than other hints and the vanilla prompt. This suggests a possible solution for employing the LLM-based generative error correction method in a code-switching scenario, where a significant mismatch exists due to low resource or domain mismatch.
TABLE X: Performance comparison of prompting LLM with the proposed linguistic hint on $\text{test}_{\texttt{man}}$ set with (the original data) and without (removing or normalizing) interjections of the SEAME dataset by employing MER (%) and utterance-level language identification accuracy (Acc %). The N-best list is decoded from the hybrid CTC/attention model with LAL
Ling. hint type $\text{test}_{\texttt{man}}$ $w/o$ $interjections$
MER $\downarrow$ Acc. $\uparrow$ MER $\downarrow$ Acc. $\uparrow$
Hybrid CTC/atten w/ LAL 16.4 - 13.6 -
no hint 16.4 - 13.7 -
LID output (LAL) 17.0 80.3 14.5 81.1
1st hypo. 16.5 93.3 13.6 95.4
hypos. vote 16.8 93.2 13.7 95.3
LAL & hypos. vote 16.8 92.9 13.7 95.3
groundtruth 15.7 - 13.4 -
TABLE XI: Performance evaluation of the proposed method on the ASRU data with different $\beta$ values in terms of utterance-level language identification accuracy (Acc. %) and MER (%)
$\beta$ for LAL MER $\downarrow$ Acc. $\uparrow$
1.0 12.1 95.5
1.5 11.9 95.7
3.0 12.0 94.1
5.0 12.2 93.4
V-F Impact of interjections
As described in Section IV-A, the SEAME dataset was collected from Singapore and Malaysia, where code-switching is more frequent than the ASRU dataset—interjections such as “lah”, “lor”, and “ya” that often occur in the former dataset. Annotating interjections is a challenging task even for individuals with bilingual expertise. Hence, interjections can introduce confusion during language modeling, leading to a degradation in the CS-ASR performance for LLMs that have not been trained on them. We therefore removed interjections from utterances in the SEAME training and $\text{test}_{\texttt{man}}$ without changing their semantic information. The CS-ASR performance and the LID accuracy of the linguistic hint prediction before and after removing interjections when incorporating LLM are shown in Table X. In addition, the first row shows the MER of the hybrid CTC/attention model with LAL. This is computed by comparing the decoded hypotheses and ground-truth transcriptions after removing interjections within them.
The results indicate that removing interjections improves the CS-ASR performance significantly due to the reduction of language confusion. However, incorporating the LLM does not benefit the CS-ASR performance except for further use of the linguistic hint with ground-truth language labels. This is consistent with results presented in Tables VIII and IX, suggesting that a correct linguistic hint can lead to performance improvement.
V-G Language identification performance
As a key component of the proposed LAL method, the LID branch often benefits from the ASR module. However, experiments in existing works have shown that LID, when employed as an auxiliary task, may not necessarily improve the ASR performance [3, 4, 51] when the LID-related parameter (i.e., $\beta$) increases.
Our primary target is to improve the CS-ASR performance of the model with the assistance of language information instead of achieving high LID performance. Although the proposed LAL generally improves the CS-ASR performance, tuning an appropriate $\beta$ for the LAL loss term in (15) is crucial for optimal performance improvement. Therefore, we present the LID performance of models trained with various $\beta$ values in Table XI.
As shown in the table, the model achieves the lowest LID performance when $\beta=5.0$. This suggests that using a higher $\beta$ value may not necessarily achieve the optimal LID performance. The highest LID performance is achieved with $\beta=1.5$. This indicates that the model optimized with an appropriate $\beta$ value can achieve high performance for both CS-ASR and LID tasks. In addition, this implies that higher ASR performance can benefit the LID branch.
VI Discussion
VI-A Frame-to-token alignment
CTC. In the hybrid CTC/attention ASR model, the frame-to-token alignment can be computed from both CTC and the cross-attention process within the ASR decoder. However, it is worth noting that CTC predictions contain $<$blank$>$ token that results in peaky behaviour [52, 53]. Since the $<$blank$>$ token lacks a language attribute and is unable to be converted into a language label, this peaky behavior leads to fewer language labels in the pseudo label sequences compared to those derived through the ASR decoder. Therefore, we use the frame-to-token alignment computed from the cross-attention attention weight matrix within the ASR decoder but not the CTC to generate pseudo language labels.
Force alignment. One can also use additional processing such as force alignment to retrieve token-level timestamps and achieve frame-to-token alignment. However, state-of-the-art multilingual ASR models, such as Whisper, require fine-tuning to achieve satisfactory performance. As presented in Table VII, the pre-trained Whisper-small model achieves MERs of 24.9%, 90.8%, and 69.7% on the ASRU and SEAME test sets, respectively. As a result, applying force alignment to code-switching speech using pre-trained multilingual ASR models is undesirable.
Results in Table VII show that the proposed LAL method outperforms performing LID with force-alignment language labels computed by the fine-tuned Whisper-small model (FA-LB). Besides, force alignment incurs a substantial computational burden, as it must be applied to the entire training dataset using fine-tuned Whisper models. While LID with FA-UB exhibits higher performance than the proposed LAL, it results in an even greater computational cost than FA-LB due to reliance on the Whisper-large model. This contradicts the motivation of our work, which is to enrich the CS-ASR model with language information without the need for an additional annotating process.
TABLE XII: Comparison of the original ASR with LAL outputs, the LM late fusion outputs, the LLM-generated outputs, and the ground-truth transcripts for the SEAME dataset, errors and corrected words are marked in red and blue, respectively.
Method Output
ASR w/ LAL 唉 呀
ah yeah close already
the yeah what happen to him ah
but 你 先 熬 一 年 先 啦
+ LM $latefusion$
ah yah close already
the yeah what happen to him ah
but 你 先 熬 一 年 先 啦
+ LLM LoRA $ft$ + ling. hint $(gt)$ ah yah
ah you are close already
the yeah what happened to him ah
but 你 现 在 熬 一 年 先 啦
Groundtruth ah yeah
ah yeah close with me
the yeah what happened to him hah
but 你 先 熬 一 年 先 啦
VI-B Where the errors happen
The errors have been analyzed to gain insights into factors affecting the ASR performance. Compared to deletion and insertion rates, a significantly higher substitution rate is observed from the aforementioned results. The high substitution rate impose a challenge in language modeling due to a larger vocabulary and language confusion arising from a code-switching scenario. Given that code-switched text may be more readily achieved than speech, developing a robust code-switching language model is desirable to address language confusion effectively.
In addition to the model performance in CS-ASR, the performance in terms of token-level language identification is worth highlighting. Compared to the CS-ASR model trained on the SEAME dataset, the model trained on the ASRU dataset can generally identify the token-level language change points. This underpins that the SEAME dataset is more challenging than the ASRU dataset since the two languages are less discriminative in the SEAME dataset.
To gain further insights into the CS-ASR performance on the SEAME dataset, we compared the decoded outputs of the ASR model incorporating LAL, and its counterparts enhanced by LM and LLM respectively. The results are juxtaposed with the ground-truth transcription in Table XII. Here, errors are marked in red and the tokens corrected by LM and LLM are marked in blue. One factor that contributes to the errors in the CS-ASR task is language confusion. For instance, the expression “ah yeah” shares both the pronunciation and semantic characteristics with its Chinese equivalent “唉 呀”. Another notable factor that contributes to errors is the liaison, where two Mandarin characters or English words can erroneously be classified as a single entity. The liaison effect can be particularly pronounced in a spontaneous code-switching speech signal.
The LLM fine-tuned on the SEAME training set demonstrated effectiveness in correcting grammatical errors. Words such as “happened” are consequently adjusted to the correct tense. However, the model is less adept at accommodating colloquial expressions. Therefore, the third and fourth cases illustrated in Table XII have been modified to a more formal expression, leading to a higher mixed error rate.
Figure 6: The visualization of multi-head cross-attention matrices within the ASR decoder layers, where the x- and y-axis denote speech frames (encoder output) and tokens, respectively.
VI-C Attention matrices
We visualize the multi-head attention matrices within various ASR decoder layers corresponding to a speech sample as shown in Fig. 6, where the x- and y-axis denote speech frames (encoder output) and tokens, respectively. The model is the best-performing CS-ASR model with the proposed LAL on the ASRU data.
Existing works have shown that shallow layers within a Transformer-based model focus on low-level speech information such as phonetics and language, while deep layers often capture high-level information such as semantic and sequential characteristics [54, 55]. This is consistent with the visualization, motivating us to employ the last decoder layer. Therefore, the average of cross-attention matrices within the last decoder is used to compute the frame-level pseudo-language labels.
Specifically, compared to the attention matrices in the third decoder layer, those in the last decoder layer capture sequential information more effectively and display a monotonic mapping of frames and tokens. However, Fig. 6(b) shows that the monotonic mapping can be blurred due to the difficulty in learning the attention pattern of each attention head during training. To mitigate this challenge, the averaged frame-to-token alignment across all attention heads is used instead of relying on a single attention head.
VII Conclusion
We proposed to align speech to languages to enhance the CS-ASR performance in both ASR and LLM-based processing. For the ASR model, we introduced a language alignment loss (LAL) to enrich the model with language information. Models equipped with the proposed LAL consistently exhibited higher CS-ASR performance than the vanilla configuration, with only a negligible increase in model parameters during training. In addition, the proposed LAL outperforms frame-level LID using language labels obtained through force alignment. These demonstrated the effectiveness of the proposed LAL. After incorporating language weights in LAL to address language imbalance in the ASRU data during training, the proposed method obtained further performance improvement. Beyond ASR, we proposed leveraging a linguistic hint, which is derived from LAL outputs and decoded hypotheses, to guide the prompting and enhance the LLM-based generative error correction. Experimental results indicate that an accurate linguistic hint can significantly improve the CS-ASR performance in scenarios involving both monolingual and code-switching utterances. Finally, the errors within the hypotheses are analyzed. The LLM fine-tuned on the SEAME data has shown to be effective in correcting grammatical errors, which, in contrast, leads to a lower ASR performance for spontaneous and colloquial speech.